# Priors in Time: Missing Inductive Biases for Language Model Interpretability

### v2

@@ 6.1 The Geometry of Stories: A Narrative-Driven Domain
Quantifying Similarity to Slow vs. Fast Moving Signals. To further quantify the straightening claim, we compute the Fourier transform of the model activations and divide the frequency spectrum into two halves at a critical frequency $f_{c}$ such that the energy (i.e., sum of squared value of phase information) in the frequencies below $f_{c}$ equals that of the remaining ones. We call the first split “slow part” of a sequence, and latter the “fast part”. We then compute the correlation matrix defined by the slow and fast parts, compute their spectrum, and analyze how similar these spectrums are to the ones defined using different SAEs’ codes. Results are shown in Fig. 6. We clearly see the spectrum extracted from the predictive component of Temporal Feature Analysis approximates that of the slow part of the representation, while the novel component’s spectrum is similar to that of the fast moving part. Meanwhile, spectra of SAEs only exhibits similarity to the fast part, suggesting they do not capture longer range dependencies necessary for interpreting narrative-driven texts like most language domains. To quantify this further, we also measure kernel similarities (CKA) between the kernels respectively defined by the slow moving and fast moving signal to the different SAE codes. Results clearly show Standard SAEs are primarily similar to fast moving part of the representation.
@@
Figure 8: Hierarchically clustering SAE codes for garden path sentences. (a) Pairwise similarity maps of the predictive code from Temporal SAE link long-distance heads and dependents that define the ultimately correct parse in garden path sentences, while standard SAEs (e.g., BatchTopK) emphasize only local, transient relations, falling for the misleading cues. (b, c) Comparing the cosine similarity of average latent code extracted from the subject phrase (SP), verb phrase (V), and object phrase (OP), we see across ambiguous garden path sentences and unambiguous control variants thereof, only the predictive component of Temporal Features Analyzers shows consistent similarity scores (as expected if the SP and V ambiguity is reflect in the latent codes).
@@ 6.2 Garden Path Sentences: Ambiguities Resolved via Temporal Structure
Garden-path sentences—e.g., The old man the boat—initially cue an incorrect local parse before a later token forces reanalysis. Language models have been shown to be able to correctly parse such ambiguous sentences, offering in fact a predictive account of human per-token surprisals (li2024incremental; hanna2024incremental; oh2023transformer). Interestingly, when using SAE codes to assess whether LLM representations offer a valid parse of the sentence, we find hierarchical clustering of SAE codes yields a parse that is suggestive of the misleading cue; meanwhile, Temporal Feature Analysis recovers the correct parse by separating the predictable, slow-moving component of the representation from the novel, fast-changing residual. Specifically, in Fig. 8, we see the predictive codes link long-distance heads and dependents that reflect the correct parse (e.g., man as verb), producing coherent similarity structure over the full span, whereas standard SAEs emphasize only transient, local changes and miss these cross-temporal constraints (see App. D.4 for more examples).
@@
Figure 9: In-context representations defined by temporal relations. (a) Schematic of the synthetic “cognitive map” task defined by park2025iclrincontextlearningrepresentations: a random walk over a small graph is concatenated into a string of tokens, which when fed into a Gemma-2-2B model leads to model next-token predictions that align with the graph structure. This behavior is exhibited in the model’s activations, since a mere 2D PCA shows the graph structure. (b) Using latent codes for low-dimension projection, we see only the predictive component of the Temporal SAE clearly recovers the graph’s geometry, whereas standard SAEs and the novel component only partially reflect it. (c) Counterfactual editing: projecting the final-token representation onto the PCA basis and replacing it with a target node’s code reliably shifts next-token predictions in a graph-consistent manner.
@@ 7.1 Geometry of latent codes in an OOD dialogue domain
We analyze a multi-turn chat domain to test whether temporal feature analysis continues to reveal slow-moving structure. We emphasize the activations in this experiment are sampled from a Gemma-2-2B-IT model, although all analyzed interpretability protocols were trained on the base (non-instruction tuned) model. Thus, in a sense, the experiments in this section also gauge the ability of different protocols to generalize out-of-distribution (OOD)—a property SAEs have been argued to partially possess when transferring between base and the corresponding instruction-tuned models (saetransfer). Specifically, we use the Ultrachat dataset, where we take the first 350 tokens from 1,000 conversations (mostly single-turn within this window, with up to four turns). For each token, we compute a shared 2D UMAP embedding from (i) raw model activations and (ii) latent codes from ReLU, TopK, and BatchTopK SAEs, as well as the novel and predictive components of the Temporal SAE. To visualize population structure clearly, we display only 100 conversations at a time to avoid clutter (Fig. 10), though fitting is applied to all 1,000.
@@ 7.2 Automatic Interpretability of Novel Codes
A central benefit of SAEs is that their latents admit automatic interpretability pipelines (“autointerp”) that attach human-readable semantics to features via activation-triggered examples, token/phrase saliency, and lightweight dashboards (e.g., GemmaScope / Neuronpedia). Our Temporal SAE adds a new, temporally aware predictive axis, but crucially it does not sacrifice these established interpretability affordances: the novel code remains compatible with standard SAE autointerp workflows. In other words, Temporal Feature Analysis augments the toolbox rather than replacing it.
To support this point, we apply a conventional autointerp loop to the novel codes: (i) collect high-activation snippets and nearest-neighbor contexts for each latent; (ii) compute direct logit attribution to inspect which vocabulary items a latent tends to support or suppress; and (iii) render compact, browsable “feature cards” that aggregate examples and scores (as in GemmaScope / Neuronpedia). While we do not perform an exhaustive causal-intervention study here, nothing in the pipeline is specific to standard SAEs; the same probes and dashboards operate unchanged on the novel codes, and we leave a systematic intervention suite to future work. Qualitative inspection confirms that the novel codes recover the kinds of monosemantic, language-level features typically reported for SAEs. Figure 11 (p. 13) shows representative examples across diverse categories—Optimism, Manipulation, University, and Slang—where the novel latents consistently activate on intuitively relevant cue phrases and contexts, and where their attribution profiles align with the expected lexical sets.
Overall, when standard SAEs produce readable feature cards, the Temporal SAE’s novel stream does as well. This complements our earlier results: the predictive stream captures slow, contextual structure (events, roles, long-range constraints), whereas the novel stream concentrates the fast, stimulus-driven information that matches inferences possible via existing SAEs, including autointerp pipelines.
@@ Acknowledgments
ESL thanks members of the Mechanisms team at Goodfire AI, particularly Tom McGrath, Owen Lewis, Jack Merullo, and Atticus Geiger for helpful discussions. ESL additionally thanks Jack Lindsey for useful comments that helped formalize the intrinsic dimensionality arguments in Sec. 4. The authors also thank David Bau for a helpful discussion about the generalization ability of SAEs. ESL and SSRH further thank Noor Sajid and members of the CRISP lab at Harvard for helpful conversations. CR is supported by a MATS extension grant. AM’s work is partially supported by the National Science Foundation (Grant No. 2530728) and Binational Science Foundation. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute at Harvard University.
@@ Appendix A Experimental details
Below, we discuss broad details of the experimental setup and measures used for corresponding evaluations.
A.1 Experiments with LLM Representations: Autocorrelation and U-Statistic
We primarily analyze Gemma-2-2B and Llama-3.1-8B models in this paper. Specifically, we precache 10K activations for the analyzed datasets for our experiments on dynamics on language model representations.
A.1.1 Autocorrelation
@@
A.1.2 U-statistic
We measure the effective dimensionality of LLM representations using a U-statistic (an unbiased estimator) based on pairwise cosine similarities. Let $\mathbf{X}_{t}=[\bm{x}_{t}^{(1)},\dots,\bm{x}_{t}^{(M)}]$ be $M$ samples of activations at time $t$ (from different timeseries). $\mathbf{G}_{t}=\mathbf{X}_{t}^{T}\mathbf{X}_{t}$ is the Gram matrix. Using the above, a U-statistic of intrinsic dimension which we employ is defined below:
$\text{U-stat}(t)=\frac{M^{2}-M}{\|\mathbf{G}_{t}\|_{F}^{2}-M}$ (5)
where $\|\mathbf{G}_{t}\|_{F}^{2}$ is the squared Frobenius norm of the Gram matrix. This quantity estimates the effective rank $1/\text{tr}(\mathbf{C}_{t}^{2})$, where $\mathbf{C}_{t}=\mathbb{E}[\bm{x}_{t}\bm{x}_{t}^{T}]$ is the second moment matrix and $\bm{x}_{t}$ is the activation vector at time $t$. Under stationarity, U-stat remains constant. When representations evolve over time, U-stat increases systematically as more orthogonal directions become active.
A.1.3 Projection Analysis
@@
A.1.4 Surrogate
@@
A.2 Analysis of Stories: UMAP Projections, Dendrograms, Similarity Heatmaps, and Event Boundary Detection
@@
A.3 Analyzing Garden Path Sentences: Dendrograms and Phrase Similarity
@@
A.4 Analyzing In-Context Representations: Visualization and Counterfactual Experiments
@@ Appendix D More Gemma Results
D.1 Hierarchical clustering of standard SAEs’ codes for story tokens
@@ Appendix F Further Results: Temporal Feature Analysis with Only Predictive Component
Appendix G Proofs
G.1 Priors on the Sparse Code for various SAEs

### v3

@@ 6.1 The Geometry of Stories: A Narrative-Driven Domain
Quantifying Similarity to Slow vs. Fast Moving Signals. To further quantify the straightening claim, we compute the temporal Fourier transform of the model activations and divide the frequency spectrum into two halves at a critical frequency $f_{c}$ such that the energy (i.e., sum of squared value of phase information) in the frequencies below $f_{c}$ equals that of the remaining ones. We call the first split “slow part” of a sequence, and latter the “fast part”. We then compute the correlation matrix defined by the slow and fast parts, compute their spectrum, and analyze how similar these spectrums are to the ones defined using different SAEs’ codes. Results are shown in Fig. 6. We clearly see the spectrum extracted from the predictive component of Temporal Feature Analysis approximates that of the slow part of the representation, while the novel component’s spectrum is similar to that of the fast moving part. Meanwhile, spectra of SAEs is primarily similar to the fast part, suggesting lack of ability to capture longer range dependencies necessary for interpreting narrative-driven texts like most language domains. To quantify this further, we also measure kernel similarities (CKA) between the kernels respectively defined by the slow moving and fast moving signal to the different SAE codes (see Tab. 4). Results clearly show Standard SAEs and novel codes are primarily similar to fast moving part of the representation, while predictive codes are similar to the slow moving part.
@@
Figure 8: Hierarchically clustering SAE codes for garden path sentences. (a) Pairwise similarity maps of the predictive code from Temporal Feature Analysis link long-distance heads and dependents that define the ultimately correct parse in garden path sentences, while standard SAEs (e.g., BatchTopK) emphasize only local, transient relations, falling for the misleading cues. (b, c) Comparing the cosine similarity of average latent code extracted from the subject phrase (SP), verb phrase (V), and object phrase (OP), we see across ambiguous garden path sentences and unambiguous control variants thereof, only the predictive component of Temporal Features Analyzers shows consistent similarity scores (as expected if the SP and V ambiguity is reflect in the latent codes).
@@ 6.2 Garden Path Sentences: Ambiguities Resolved via Temporal Structure
Garden-path sentences—e.g., The old man the boat—initially cue an incorrect local parse before a later token forces reanalysis. Language models have been shown to be able to correctly parse such ambiguous sentences, offering in fact a predictive account of human per-token surprisals (li2024incremental; hanna2024incremental; oh2023transformer). Interestingly, when using SAE codes to assess whether LM activations offer a valid parse of the sentence, we find hierarchical clustering of SAE codes yields a parse that is suggestive of the misleading cue; meanwhile, Temporal Feature Analysis recovers the correct parse by separating the predictable, slow-moving component of the representation from the novel, fast-changing residual. Specifically, in Fig. 8, we see the predictive codes link long-distance heads and dependents that reflect the correct parse (e.g., man as verb), producing coherent similarity structure over the full span, whereas standard SAEs emphasize only transient, local changes and miss these cross-temporal constraints (see App. D.4 for more examples).
@@
Figure 9: In-context representations defined by temporal relations. (a) Schematic of the synthetic “cognitive map” task defined by park2025iclrincontextlearningrepresentations: a random walk over a small graph is concatenated into a string of tokens, which when fed into a Gemma-2-2B model leads to model next-token predictions that align with the graph structure. This behavior is exhibited in the model’s activations, since a mere 2D PCA shows the graph structure. (b) Using latent codes for low-dimension projection, we see only the predictive component of Temporal Feature Analysis recovers the graph’s geometry, whereas standard SAEs and the novel component only partially reflect it. (c) Counterfactual editing: projecting the final-token representation onto the PCA basis and replacing it with a target node’s code reliably shifts next-token predictions in a graph-consistent manner.
@@ 7.1 Geometry of latent codes in an OOD dialogue domain
We analyze a multi-turn chat domain to test whether temporal feature analysis continues to reveal slow-moving structure. We emphasize the activations in this experiment are sampled from a Gemma-2-2B-IT model, although all analyzed interpretability protocols were trained on the base (non-instruction tuned) model. Thus, in a sense, the experiments in this section also gauge the ability of different protocols to generalize out-of-distribution (OOD)—a property SAEs have been argued to partially possess when transferring between base and the corresponding instruction-tuned models (saetransfer). Specifically, we use the Ultrachat dataset, where we take the first 350 tokens from 1,000 conversations (mostly single-turn within this window, with up to four turns). For each token, we compute a shared 2D UMAP embedding from (i) raw model activations and (ii) latent codes from ReLU, TopK, and BatchTopK SAEs, as well as the novel and predictive components of Temporal Feature Analyzers. To visualize population structure clearly, we display only 100 conversations at a time to avoid clutter (Fig. 10), though fitting is applied to all 1,000.
@@ 7.2 Automatic Interpretability of Novel Codes
A central benefit of SAEs is that their latents admit automatic interpretability pipelines (“autointerp”) that attach human-readable semantics to features via activation-triggered examples, token/phrase saliency, and lightweight dashboards (e.g., GemmaScope / Neuronpedia). Our Temporal Feature Analyzer adds a new, temporally aware predictive axis, but crucially it does not sacrifice these established interpretability affordances: the novel code remains compatible with standard SAE autointerp workflows. In other words, Temporal Feature Analysis augments the toolbox rather than replacing it. To support this point, we apply a conventional autointerp loop to the novel codes: (i) collect high-activation snippets and nearest-neighbor contexts for each latent; (ii) compute direct logit attribution to inspect which vocabulary items a latent tends to support or suppress; and (iii) render compact, browsable “feature cards” that aggregate examples and scores (as in GemmaScope / Neuronpedia). While we do not perform an exhaustive causal-intervention study here, nothing in the pipeline is specific to standard SAEs; the same probes and dashboards operate unchanged on the novel codes, and we leave a systematic intervention suite to future work. Qualitative inspection confirms that the novel codes recover the kinds of monosemantic, language-level features typically reported for SAEs. Figure 11 (p. 13) shows representative examples across diverse categories—Optimism, Manipulation, University, and Slang—where the novel latents consistently activate on intuitively relevant cue phrases and contexts, and where their attribution profiles align with the expected lexical sets. Overall, when standard SAEs produce readable feature cards, the Temporal Feature Analysis’s novel stream does as well. This complements our earlier results: the predictive stream captures slow, contextual structure (events, roles, long-range constraints), whereas the novel stream concentrates the fast, stimulus-driven information that matches inferences possible via existing SAEs, including autointerp pipelines.
@@ Acknowledgments
ESL thanks members of the Mechanisms team at Goodfire AI, particularly Tom McGrath, Owen Lewis, Jack Merullo, and Atticus Geiger for helpful discussions. ESL additionally thanks Jack Lindsey for useful comments that helped formalize the intrinsic dimensionality arguments in Sec. 4, and Raphael Sarfati for suggesting the Tortuosity metric used in Fig. 5. The authors also thank David Bau for a helpful discussion about the generalization ability of SAEs, and David Klindt for several useful comments on an earlier draft of the paper. ESL and SSRH further thank Noor Sajid and members of the CRISP lab at Harvard for helpful conversations. CR is supported by a MATS extension grant. AM’s work is partially supported by the National Science Foundation (Grant No. 2530728) and Binational Science Foundation. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute at Harvard University.
@@ Appendix A Experimental details
Below, we discuss broad details of the experimental setup and measures used for corresponding evaluations. For experiments with LM representations, we primarily analyze Gemma-2-2B and Llama-3.1-8B models in this paper. Specifically, we precache 10K activations for the analyzed datasets for our experiments on dynamics on language model representations.
A.1 Autocorrelation Between Activations
@@
A.2 Projection Analysis
@@
A.3 U-statistic
We initially experimented with more well-known measures of intrinsic dimensionality, but found them to either be too computationally expensive [costa2005estimating] or saturate due to inappropriate assumptions (e.g., data satisfying a Euclidean geometry) [carter2009local]. We thus defined an intrinsic dimensionality measure for a specific generative process for activations that is motivated by the linear representation hypothesis [park2023linear, elhage2022superposition, costa2025flat, arora2018linear]. Specifically, assume that each observation (activation) $x_{t}\in\mathbb{R}^{D}$ is a unit-norm vector generated from an (unknown) overcomplete dictionary $V=[v_{1},\dots,v_{N}]\in\mathbb{R}^{D\times N},$ where $D<N$, whose columns are approximately orthonormal, in the sense that $\langle v_{i},v_{j}\rangle$ is small for $i\neq j$ and $\|v_{i}\|_{2}=1$. At time $t$, the observation is $x_{t}=Va_{t}$ and $\|x_{t}\|_{2}=1$, where $a_{t}\in\mathbb{R}^{N}$ is a random coefficient vector with a time-dependent distribution $P_{t}(a)$. Thus the process is nonstationary through the evolution of $P_{t}$ (and hence $a_{t}$), while the dictionary $V$ is fixed, matching the phenomenology observed in Sec. 3.
Under the assumptions above, we measure the intrinsic dimensionality of LM activations based on pairwise cosine similarities. Specifically, let $\mathbf{X}_{t}=[\bm{x}_{t}^{(1)},\dots,\bm{x}_{t}^{(M)}]$ be $M$ samples of normalized activations at time $t$ (from different timeseries). Define the gram matrix $\mathbf{G}_{t}=\mathbf{X}_{t}^{T}\mathbf{X}_{t}$. Then, a U-statistic of intrinsic dimension can be defined as follows:
$\text{U-stat}(t)=\frac{M^{2}-M}{\|\mathbf{G}_{t}\|_{F}^{2}-M},$ (7)
where $\|\mathbf{G}_{t}\|_{F}^{2}$ is the squared Frobenius norm of the Gram matrix. As we show in the following subsection, this quantity is an estimator of the effective rank $\nicefrac{{1}}{{\text{tr}(\mathbf{C}_{t}^{2})}}$, where $\mathbf{C}_{t}=\mathbb{E}[\bm{x}_{t}\bm{x}_{t}^{T}]$ is the second moment matrix and $\bm{x}_{t}$ is the activation vector at time $t$. Under stationarity, U-stat remains constant; when representations evolve over time, U-stat increases systematically as more orthogonal directions become active.
A.3.1 Why is U-stat an intrinsic dimensionality measure
Let $P_{t}$ denote the distribution of $x_{t}$ at time $t$ and define the (time-local) second-moment matrix $C_{t}\;:=\;\mathbb{E}_{x\sim P_{t}}[xx^{\top}]\in\mathbb{R}^{D\times D}$. Assuming $\|x\|_{2}=1$, we have $\operatorname{tr}(C_{t})=1$. Given $M$ samples $\{x_{t,1},\dots,x_{t,M}\}$ drawn from $P_{t}$ (e.g., from a time window or an ensemble of trajectories at time $t$), we form the Gram matrix $G_{t}\in\mathbb{R}^{M\times M}$ with entries $(G_{t})_{ij}=x_{t,i}^{\top}x_{t,j}$. The corresponding U-statistic
$\hat{U}_{t}\;:=\;\frac{1}{M(M-1)}\sum_{i\neq j}(x_{t,i}^{\top}x_{t,j})^{2}\;=\;\frac{\|G_{t}\|_{F}^{2}-M}{M^{2}-M}$
is a consistent estimator of
$U_{t}\;:=\;\mathbb{E}_{x,x^{\prime}\sim P_{t}}[(x^{\top}x^{\prime})^{2}]\;=\;\operatorname{tr}(C_{t}^{2}),$
where $x$ and $x^{\prime}$ are independent draws from $P_{t}$. Thus $\hat{U}_{t}$ provides a basis-free estimate of $\operatorname{tr}(C_{t}^{2})$.
Now, a natural, rotation-invariant notion of intrinsic dimensionality at time $t$ is the (order-2) effective rank of $C_{t}$, $d_{\mathrm{eff}}(t)\;:=\;\frac{(\operatorname{tr}C_{t})^{2}}{\operatorname{tr}(C_{t}^{2})}$. For unit-norm data $\operatorname{tr}(C_{t})=1$, so
$d_{\mathrm{eff}}(t)\;=\;\frac{1}{\operatorname{tr}(C_{t}^{2})}\;=\;\frac{1}{U_{t}},\qquad\hat{d}_{\mathrm{eff}}(t)\;:=\;\frac{1}{\hat{U}_{t}}\;=\;\frac{M^{2}-M}{\|G_{t}\|_{F}^{2}-M}.$
Let $\lambda_{1}(t),\dots,\lambda_{D}(t)$ be the eigenvalues of $C_{t}$. Then $\operatorname{tr}(C_{t}^{2})=\sum_{i=1}^{D}\lambda_{i}(t)^{2}$. This yields the following.
$d_{\mathrm{eff}}(t)=\frac{1}{\sum_{i=1}^{D}\lambda_{i}(t)^{2}}.$
Now if the process at time $t$ spreads its energy equally over exactly $k$ orthogonal directions, then $\lambda_{1}(t)=\dots=\lambda_{k}(t)=1/k$ and $\lambda_{k+1}(t)=\dots=0$, and, hence, $d_{\mathrm{eff}}(t)=k$. However, if most of the energy lies in a single direction, then $\sum_{i}\lambda_{i}(t)^{2}$ is close to $1$ and $d_{\mathrm{eff}}(t)$ is close to $1$. Thus, $d_{\mathrm{eff}}(t)$ behaves like the “number of active directions” in which the process varies at time $t$. This is especially appropriate as a measure for our assumed generative process, i.e., $x_{t}=Va_{t}$, which yields
$C_{t}=\mathbb{E}[x_{t}x_{t}^{\top}]=\mathbb{E}[Va_{t}a_{t}^{\top}V^{\top}]=VS_{t}V^{\top},\qquad S_{t}:=\mathbb{E}[a_{t}a_{t}^{\top}].$
If $V$ is close to orthonormal and $S_{t}$ is close to diagonal, with
$e_{i}(t):=\mathbb{E}[a_{t,i}^{2}]$
denoting the average energy of coefficient $i$, then
$\operatorname{tr}(C_{t}^{2})\approx\sum_{i=1}^{N}e_{i}(t)^{2},\qquad\sum_{i=1}^{N}e_{i}(t)\approx 1.$
Hence
$d_{\mathrm{eff}}(t)=\frac{1}{\operatorname{tr}(C_{t}^{2})}\approx\frac{1}{\sum_{i=1}^{N}e_{i}(t)^{2}},$
which is the usual “effective number” of coefficients that carry energy at time $t$. In particular, if the energy is spread evenly across $k$ dictionary atoms, then $d_{\mathrm{eff}}(t)\approx k$. Therefore $d_{\mathrm{eff}}(t)$, estimated from the Gram U-statistic, is a simple and natural measure of intrinsic dimensionality for this generative process.
A.4 Surrogate
@@
A.5 Analysis of Stories: UMAP Projections, Dendrograms, Similarity Heatmaps, and Event Boundary Detection
@@
A.6 Analyzing Garden Path Sentences: Dendrograms and Phrase Similarity
@@
A.7 Analyzing In-Context Representations: Visualization and Counterfactual Experiments
@@ Appendix D More Gemma Results
D.1 Hierarchical clustering of codes from story tokens
@@ Appendix F Further Results: Temporal Feature Analysis with Only Predictive Component
Appendix G Further Results: Emergent Separation of Predictive and Novel Code in Temporal Feature Analysis
Figure 58: Separation of Predictive and Novel Dictionaries. Predictive and novel codes for sequences drawn from three datasets—SimpleStories (top), Webtext (middle), and Code (bottom)—are binarized to zero (black) and non-zero (white). The dictionary elements are sorted by the sum of non-zero activations in the predictive code across the sequence. The ordering obtained from the predictive code is also applied to the novel code. The red dashed line marks the separation of 90% of non-zero activation counts; only 10% are overlapping into the other subset. Additionally, the effective rank of the predictive codes is two orders of magnitudes lower than the average number of non-zero dictionary elements (Mean L0).
The Temporal Feature Analyzers studied in the main paper share a dictionary for both predictive and novel codes. We now investigate if there is any shared structure in the two dictionaries. In particular, given the predictive and novel codes play a different role, it is plausible the SAE learns to approximately split the dictionary into two parts: one responsible for computing the predictive code and the other for novel one. Specifically, we investigate a Temporal Feature Analyzer trained on Gemma-2-2B Layer 12 residual stream activations. Results are shown in Fig. 58 and show a separation of dictionary elements for sequences drawn from three datasets: SimpleStories, Webtext, and Code. Specifically, we find $\sim$2K dictionary elements participate in defining the predictive code, while the remaining $\sim$8K primarily participate in defining the novel code. However, the absolute count is merely indicative of the L0 sparsity of a code, which may not reflect how many directions are actually used to define the code—estimating latter requires computing the rank of the code (in fact, we note that rank-sparsity is a well-known alternative to L0 sparsity in dictionary learning literature [elad2010sparse]). As we show in Fig. 58, the effective rank of the predictive codes ($\sim$20–30) is substantially lower than both the effective rank ($\sim$390–410) and absolute L0 sparsity of novel codes.
Overall, the posthoc analysis above shows that predictive and novel codes largely use separate subsets of dictionary elements. This emergent disentanglement of the two components motivates one to preemptively split the dictionary into two components—one responsible for the predictive code and other for novel one. Our preliminary experiments show a Temporal SAE trained with this split-dictionary architecture is similarly performant as the tied dictionary one, resulting in predictive / novel codes with effective ranks $\sim$20–30 / $\sim$390–420, and a slightly better overall loss. It is possible optimizing hyperparameters (e.g., having different expansion factors for the two components) can make this architecture more performant than the tied dictionary one, but we leave a further characterization to future work.
Appendix H Proofs
H.1 Priors on the Sparse Code for various SAEs